# Scene Graph Generation in Large-Size VHR Satellite Imagery: A Large-Scale Dataset and A Context-Aware Approach

### v2

@@ 5.1 Benchmark
Task Settings. For the OBD task, we establish two types of OBD benchmarks: HBB-based and OBB-based detectors in large-size VHR SAI. For the SGG task, we evaluate SGG in large-size VHR SAI following three conventional sub-tasks of SGG in natural imagery.
Predicate Classification (PredCls). Given the object categories and their bounding boxes, it is necessary to predict pairs containing rich knowledge and relationship types of these pairs in large-size VHR SAI. This sub-task is to verify the performance of the model in pair pruning and relationship prediction without other interfering factors.
Scene Graph Classification (SGCls). Given only the bounding boxes of objects, it is necessary to recognize the categories of both subjects and objects, and thus predict high-value pairs as well as relationship types of these pairs in large-size VHR SAI.
Scene Graph Generation (SGGens). Given a large-size VHR SAI, it is direct to predict high-value triplets $<$subject, relationship, object$>$. These triplets encompass the bounding boxes and categories of both subjects and objects, as well as the categories of their relationships.
@@
To evaluate the performance of the SGG model more comprehensively, the harmonized mean (HMR@K) of both MR@K and mMR@K is taken as the comprehensive evaluation metric for the SGG task in large-size VHR SAI. For all tasks, the prediction can be considered as a true positive (TP) result only when it matches the ground-truth triplet $<$subject, relationship, object$>$ and has at least 0.5 IoU between the bounding boxes of subject-object and the bounding boxes of ground-truth. As for the K, instead of the setting (K= 50/100), which is commonly used by the SGG task in natural imagery, we adopt the setting (K= 1500/2000). This is because, in large-size VHR SAI, the average number of triplets per image far exceeds that in natural imagery.
@@ 5.2 Implementation details
Object Detector. Taking the number of instances and the type of scenarios as the basis of division, we split the dataset into train set (60%), val set (20%), and test set (20%). The detectors are trained on RSG train set using AdamW [100] as an default optimizer, and the performance is reported on RSG test set. All the listed models are trained on NVIDIA RTX3090 GPUs. We set the batch size to 2 and the initial learning rate to $1\times 10^{-4}$. The mean average precision (mAP) is calculated following PASCAL VOC 07 [101].
Scene Graph Generation. On top of the frozen detectors, we train all SGG models on RSG train set using SGD as an optimizer and evaluate them on RSG test set. Batch size and initial learning rate are 4 and $1\times 10^{-3}$ for the PredCls and SGCls tasks, and 2 and $1\times 10^{-3}$ for the SGDet task, respectively. For the SGDet task, object detection is performed using a Per-Class NMS [45, 43] with 0.5 IoU. Considering the computational resource consumption caused by numerous invalid object pairs in large-size VHR SAI, we use the proposed PPG network in the overall framework, which enables the SGG task in large-size VHR SAI to be realized in common computational conditions (e.g., one single GPU).
Figure 9: Qualitative comparisons between proposed RPCM and other methods on different scenarios. Blue edges represent the ground-truth relationships correctly predicted, purple edges indicate ground-truth relationships that failed to be detected, and green edges are reasonable relationships predicted by the model but not annotated in the ground-truth.
TABLE IV: Baseline results (%) of SGG on RSG test set. Specifically, our model outperforms previous state-of-the-art models by an absolute 5.17%/4.76% on HMR@1500/2000 for the SGCls task and 1.27%/1.82% on HMR@1500/2000 for the SGDet task on the OBB-based detector.
@@
OBD Type Relationship Prediction Venue MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000
@@
Figure 10: The visualization results of pair score matrix. Blue edges represent the ground-truth pairs correctly predicted, purple edges indicate ground-truth pairs that failed to be detected, and green edges are reasonable pairs predicted by the model but not annotated in the ground-truth.
TABLE V: Comparison of our PPG and other pair pruning methods on RSG test set for the PredCls, SGCls, and SGDet tasks.
@@
5.3 Results and Analysis for OBD
Comparison with Baselines. To explore the properties of RSG and provide guidelines for future OBD task in large-size VHR SAI, we conduct a comprehensive evaluation of about 30 OBD methods and analyze the results. Table II provides a comprehensive benchmark on RSG, including single/two-stage, CNN/Transformer, anchor-based/anchor-free and supervised/weakly-supervised methods. We adopt MMDetection [^1] https://github.com/Zhuzi24/RSG-MMDetection [102] and MMRotate [^2] https://github.com/yangxue0827/RSG-MMRotate [103] as the toolkits and all experiments are based on the standard ‘1x’ (12 epochs) training schedule. It can be observed that our HOD-Net significantly outperforms other methods under the same backbone (ResNet50 [73]) for both HBB-based and OBB-based detectors. To provide more accurate OBD results for the SGG task in large-size VHR SAI, we evaluate three OBD strategies on HBB-based and OBB-based detectors (see Table III): resizing, cropping, and our HOD-Net. It can be seen that our HOD-Net achieves 53.2% and 55.9% mAP on the HBB-based and OBB-based detectors, respectively (using a 0.5 IoU threshold). Specifically, under the same backbone (Swin-L [72]), it achieves a 10.6%/19.5% improvement in mAP compared to the baseline Faster R-CNN detector (cropping/resizing) in the HBB detection task, and a 12.9%/23.9% improvement in mAP metric compared to the baseline Oriented R-CNN detector (cropping/resizing) in the OBB detection task. Notably, for objects with extreme aspect ratios (e.g., runways, ship locks), our HOD-Net outperforms other methods in OBB detection by a great margin of (34.9%/31.2%). Above experiments indicate that HOD-Net can provide more comprehensive and accurate object representation for SGG in large-size VHR SAI.
Qualitative Results and Visualization. In Fig. 8, we further give the visualizations of different strategies on HBB-based detectors and OBB-based detectors. It can be seen that many large objects (e.g., docks, aprons) cannot be detected holistically due to the integrity being destroyed in the detection based on cropping strategy, and a lot of small objects (e.g., airplanes, boats, cranes, boarding bridges) cannot be detected correctly due to the loss of information in the detection based on resizing strategy. The proposed HOD-Net shows superiority in the detection results of multi-scale objects and achieves almost the same results as the annotation for objects with extreme aspect ratios (e.g., runways). The above results show that HOD-Net can provide mandatory support for the SGG task in large-size VHR SAI.
@@ 5.4 Results and Analysis for SGG
Figure 11: Visualization of different PBA iters L. Blue edges represent the correctly predicted relationships, purple edges indicate incorrectly predicted relationships, and green edges are reasonable relationships predicted by the model but not annotated in the ground-truth.
Qualitative Results and Visualization. The SGG visualization results of different models on the RSG dataset are shown in Fig. 9. It can be seen that HETSGG has a limited ability to learn discriminative details(e.g., HETSGG fails to discriminate specific relationships $<$ship, away from/approach, dock$>$ by the wake trailing of the ship). PE-Net does not have the ability to constrain relationship prediction by relying on contextual reasoning (e.g., inferring the unreasonable triplet $<$airplane0, parking in the same apron with, airplane2$>$ via triplet $<$airplane2, parallelly parked on, apron2$>$ and triplet $<$airplane0, parallelly parked on, apron1$>$). In summary, our RPCM network can effectively guide the model in learning the discriminative details of different relationships, demonstrating a strong capacity for context learning and inference.
5.5 Ablation Study on SGG
5.5.1 Effect of PPG Network
Table V gives the quantitative comparison of the proposed PPG network with other pair pruning methods on the RSG dataset when the sorted top 10,000 pairs are selected. It can be seen that the ABBS method based on statistics has only a weak improvement over the randomized method in all metrics, which shows that the semantic and spatial combination types of the object pairs in large-size VHR SAI are so complex that it is difficult to obtain combination patterns of objects by statistics. The score matrix, as depicted in the first row of Fig. 10, becomes sparse after the PPG network learning, effectively retaining the annotated pairs. This highlights the robust learning capacity of the proposed PPG network when dealing with pairs containing rich knowledge. The local visualization shows that the PPG network can retain the annotated pairs with high scores, as well as some pairs that are non-annotated but may contain high-value relationships, this is because the annotation of triplets in the real annotation is hardly exhaustive, so achieving the same sparsity as ground-truth is not necessary.
TABLE VI: Comparison with different PBA iters L on MR@1500/2000, mMR@1500/2000, and HMR@1500/2000.
@@ 5.5.1 Effect of PPG Network
5.5.2 Effect of PBA Iters
In Table VI, we compare the results on PredCls for different numbers of PBA iters L (L is set to 1, 2, 3, 4, 5). We find that the number of iters has little effect on the results and our model is robust to the number of iters. Notably, the HMR@K keeps improving with the increase in iters when iters are less than 4. This suggests that proper contextual information can effectively improve the model performance, but excessive introduction of contextual information may cause disturbance. In this paper, the number of iters for PBA is set to 4.
5.5.3 Effect of Different Components
To investigate the effectiveness of the components of the RPCM, we perform component ablation experiments by gradually removing them from the RPCM on the PredCls task, as shown in Table VII. Not surprisingly, the experimental results show that all three components, object context augmentation, relationship context augmentation, and prototype matching, are important for the performance of SGG. Specifically, prototype matching enables the model to capture the discriminative features of different relationship categories in the semantic space, which in turn effectively distinguishes different categories with strong similarities. The object augmentation and relationship augmentation modules assist the model in learning contextual reasoning constraints for relationship prediction, and the visualization in Fig. 9 also reflects it well.
@@ 6 Conclusion
In this paper, we propose a large-scale dataset named RSG for SGG in large-size VHR SAI. The RSG dataset covers 1,273 complex scenarios worldwide, with image sizes ranging from 512 × 768 to 27,860 × 31,096 pixels, including more than 210,000 objects (with both OBB and HBB annotations) and more than 400,000 triplet annotations. The large image size, the extensive sample volume, and the diversity of object scale and relationship semantics make RSG a valuable dataset, providing indispensable data support for advancing a new challenging and meaningful task: SGG in large-size VHR SAI. Furthermore, we propose the CAC framework, a solution tailored for the SGG task in large-size VHR SAI, to achieve a cascading understanding of SAI from three levels: object detection, pair pruning and relationship prediction. To facilitate the development of SGG in large VHR SAI, this paper releases a SAI-oriented toolkit containing various SGG methods and develops a benchmark in which the effectiveness of the proposed CAC framework is empirically validated. In future work, we will continue to enrich the RSG dataset in terms of its sample volume, object categories, and relationship categories. Additionally, we will strive to explore methods that can simultaneously improve the accuracy of OBD and SGG in large-size VHR SAI, thereby maximizing the potential of SGG in various downstream tasks of SAI, such as image retrieval, image captioning and visual question answering.
@@ Acknowledgment
We sincerely thank the anonymous editors and reviewers for their insightful comments and suggestions. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.

### v3

@@ 5.1 Benchmark
Task Settings. For the OBD task, we establish two types of OBD benchmarks: HBB-based and OBB-based detectors in large-size VHR SAI. For the SGG task, we evaluate SGG in large-size VHR SAI following three conventional sub-tasks of SGG in natural imagery: Predicate Classification (PredCls), Scene Graph Classification (SGCls) and Scene Graph Detection (SGDet).
@@
To evaluate the performance of the SGG model more comprehensively, the harmonized mean (HMR@K) of both MR@K and mMR@K is taken as the comprehensive evaluation metric for the SGG task in large-size VHR SAI. For all tasks, the prediction can be considered as a true positive (TP) result only when it matches the ground-truth triplet $<$subject, relationship, object$>$ and has at least 0.5 IoU between the bounding boxes of subject-object and the bounding boxes of ground-truth. As for the K, instead of the setting (K= 50/100), which is commonly used by the SGG task in natural imagery, we adopt the setting (K= 1,500/2,000) due to the average number of triplets per large-size VHR SAI far exceeds that in natural imagery.
@@ 5.2 Implementation details
Object Detector. Taking the number of instances and the type of scenarios as the basis of division, we split the dataset into train set (60%), val set (20%), and test set (20%). The detectors are trained on STAR train set using AdamW [95] as an default optimizer, and the performance is reported on STAR test set. All the listed models are trained on NVIDIA RTX3090 GPUs. We set the batch size to 2 and the initial learning rate to $1\times 10^{-4}$. The mean average precision (mAP) is calculated following PASCAL VOC 07 [96].
Scene Graph Generation. On top of the frozen detectors, we train all SGG models on STAR train set using SGD as an optimizer and evaluate them on STAR test set. Batch size and initial learning rate are 4 and $1\times 10^{-3}$ for the PredCls and SGCls tasks, and 2 and $1\times 10^{-3}$ for the SGDet task, respectively. For the SGDet task, object detection is performed using a Per-Class NMS [43, 41] with 0.5 IoU. Considering the computational resource consumption caused by numerous invalid object pairs in large-size VHR SAI, we use the proposed PPG network in the overall framework, which enables the SGG task in large-size VHR SAI to be realized in common computational conditions (e.g., one single GPU).
Figure 9: The visualization results of the SGG task. Blue edges represent the correctly predicted relationships, purple edges indicate the ground-truth (GT) relationships that failed to be detected, and green edges are reasonable relationships predicted by the model but not annotated in the GT.
TABLE IV: Baseline results (%) of SGG on STAR test set for the PredCls, SGCls and SGDet tasks.
@@
OBD
Type
Relationship
Prediction
Venue MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000 MR@1500/2000 mMR@1500/2000 HMR@1500/2000
@@
5.3 Results and Analysis for OBD
Comparison with Baselines. To explore the properties of STAR and provide guidelines for future OBD task in large-size VHR SAI, we conduct a comprehensive evaluation of about 30 OBD methods and analyze the results. Table II provides a comprehensive benchmark on STAR, including single/two-stage, CNN/Transformer, anchor-based/anchor-free and supervised/weakly-supervised methods. We adopt MMDetection [^1] https://github.com/Zhuzi24/STAR-MMDetection [97] and MMRotate [^2] https://github.com/yangxue0827/STAR-MMRotate [98] as the toolkits and all experiments are based on the standard ‘1x’ (12 epochs) training schedule. It can be observed that our HOD-Net significantly outperforms other methods under the same backbone (ResNet50 [68]) for both HBB-based and OBB-based detectors. To provide more accurate OBD results for the SGG task in large-size VHR SAI, we evaluate three OBD strategies on HBB-based and OBB-based detectors (see Table III): resizing, cropping, and our HOD-Net. It can be seen that our HOD-Net achieves 53.2% and 55.9% mAP on the HBB-based and OBB-based detectors, respectively (using a 0.5 IoU threshold). Specifically, under the same backbone (Swin-L [67]), it achieves a 10.6%/19.5% improvement in mAP compared to the baseline Faster R-CNN detector (cropping/resizing) in the HBB detection task, and a 12.9%/23.9% improvement in mAP metric compared to the baseline Oriented R-CNN detector (cropping/resizing) in the OBB detection task. Notably, for objects with extreme aspect ratios (e.g., runways, ship locks), our HOD-Net outperforms other methods in OBB detection by a great margin of (34.9%/31.2%). Above experiments indicate that HOD-Net can provide more comprehensive and accurate object representation for SGG in large-size VHR SAI.
Qualitative Results and Visualization. In Fig. 8, we further give the visualizations of different methods on OBB-based detectors and OBB-based detectors. It can be seen that many large objects (e.g., docks, aprons) cannot be detected holistically due to the integrity being destroyed in the detection based on cropping strategy, and a lot of small objects (e.g., airplanes, boats, cranes, boarding bridges) cannot be detected correctly due to the loss of information in the detection based on resizing strategy. The proposed HOD-Net shows superiority in the detection results of multi-scale objects and achieves almost the same results as the annotation for objects with extreme aspect ratios (e.g., runways). The above results show that HOD-Net can provide mandatory support for the SGG task in large-size VHR SAI.
Figure 10: The visualization results of pair score matrix. Blue edges represent the ground-truth (GT) pairs correctly predicted, purple edges indicate the GT pairs that failed to be detected, and green edges are reasonable pairs predicted by the model but not annotated in the GT.
TABLE V: Comparison of our PPG and other pair pruning methods on STAR test set for the PredCls, SGCls, and SGDet tasks.
@@ 5.4 Results and Analysis for SGG
Qualitative Results and Visualization. The SGG visualization results of different models on the STAR dataset are shown in Fig. 9. HETSGG has a limited ability to learn discriminative details(e.g., HETSGG fails to discriminate specific relationships $<$ship, away from/approach, dock$>$ by the wake trailing of the ship). PE-Net does not have the ability to constrain relationship prediction by relying on contextual reasoning (e.g., inferring the unreasonable triplet $<$airplane0, parking in the same apron with, airplane2$>$ via triplet $<$airplane2, parallelly parked on, apron2$>$ and triplet $<$airplane0, parallelly parked on, apron1$>$). In summary, our RPCM network can effectively guide the model in learning the discriminative details of different relationships.
TABLE VI: Comparison with different PBA iters L.
@@
5.5 Ablation Study on SGG
5.5.1 Effect of PPG Network
Table V gives the quantitative comparison of the proposed PPG network with other pair pruning methods on the STAR dataset when the sorted top 10,000 pairs are selected. It can be seen that the ABBS method based on statistics has only a weak improvement over the randomized method in all metrics, which shows that the semantic and spatial combination types of the object pairs in large-size VHR SAI are so complex that it is difficult to obtain combination patterns of objects by statistics. The score matrix, as depicted in the first row of Fig. 10, becomes sparse after the PPG network learning, effectively retaining the annotated pairs. This highlights the robust learning capacity of the proposed PPG network when dealing with pairs containing rich knowledge. The local visualization shows that the PPG network can retain the annotated pairs with high scores, as well as some pairs that are non-annotated but may contain high-value relationships, this is because the annotation of triplets in the real annotation is hardly exhaustive, so achieving the same sparsity as ground-truth is not necessary.
@@ 5.5.1 Effect of PPG Network
Figure 11: Visualization of different PBA iters L. Blue edges represent the correctly predicted relationships, purple edges indicate incorrectly predicted relationships, and green edges are reasonable relationships predicted by the model but not annotated in the ground-truth (GT).
5.5.2 Effect of PBA Iterations
In Table VI, we compare the results on PredCls for different numbers of PBA iters L (L is set to 1, 2, 3, 4, 5). We find that the number of iters has little effect on our model. Notably, the HMR@K keeps improving with the increase in iters when iters are less than 4. This suggests that proper contextual information can effectively improve the model performance, but excessive introduction of contextual information may cause disturbance. In this paper, the number of iters for PBA is set to 4.
5.5.3 Effect of Different Components
We perform component ablation by gradually removing them from the RPCM on the PredCls task, as shown in Table VII. Not surprisingly, the experimental results show that all three components, object context augmentation, relationship context augmentation, and prototype matching, are important for the performance of SGG. Specifically, prototype matching enables it to capture the discriminative features of different relationship categories in the semantic space, which in turn effectively distinguishes different categories with strong similarities. The object augmentation and relationship augmentation modules assist the model in learning contextual reasoning constraints for relationship prediction, as shown by the visualization in Fig. 9.
@@ 6 Conclusion
We have proposed a first-ever dataset named STAR for SGG in large-size VHR SAI. It covers 1,273 complex scenarios worldwide, with image sizes ranging from 512 × 768 to 27,860 × 31,096 pixels, including more than 210,000 objects (with both OBB and HBB annotations) and more than 400,000 triplet annotations. The large image size, the extensive sample volume, and the diversity of object scale and relationship semantics make STAR a valuable dataset, providing indispensable data support for advancing a new challenging and meaningful task: SGG in large-size VHR SAI. Furthermore, we propose the CAC framework, for the SGG task in large-size VHR SAI, to achieve a cascading understanding of SAI from three levels: object detection, pair pruning and relationship prediction. To facilitate the development of SGG in large VHR SAI, this paper releases an open-source toolkit containing various SGG methods and develops a large-scale benchmark in which the effectiveness of the proposed CAC framework is empirically validated.
@@ Acknowledgment
This work was partly supported by the National Natural Science Foundation of China (42371321 and 42030102) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).